3. A 20-g bullet flying horizontally becomes embedded in a 4-kg wooden block suspended from the ceiling. What was the speed of the bullet if the block swung 5 cm up from its initial level above ground?
Answer #1

When the block swung its : kinetic energy that it gained during collision gets converted completely into gravitational potential energy at height of 5cm or 0.05 m , so applying conservation of energy after bullet get embedded in block

\frac{mv^2}{2 } = mgh

m = combined mass of block and bullet

g = 9.8 m/s2

h = 0.05 m

v = speed of bullet and block combined just after collision

\frac{v^2}{2 } = gh

v^2 = 2gh

v= \sqrt{2gh}

v= \sqrt{2*9.8*0.05}

v\approx 1 m/s

In horizontal direction there is no external force on system during collision so now applying conservation of momentum in horizontal direction during collision

momentum of system before collison = mass of bullet * speed of bullet = 0.02 * vb

momentum of system after collision = mv = (4+0.02) * 1 = 4.02 kg m/s

equating both the momentums

0.02 * vb = 4.02

vb = 201 m/s